Two trains are moving with equal speed in opposite direction along to railway track if the wind is blowing with the speed u along the track so that relative velocity of the train with respect to the wind are in the ratio 1 is to 2 then the speed of each train must be

Grade:11

1 Answers

Arun
5 years ago
Since, two paralell trains are moving in the speed, they are having a uniform speed. 
However, the velocities are in he ratio, 2:1.
This means that the train that is moving along with the speed of of wind = (v+u) and the train moving against the speed of the wind = (v-u). 
Therefore, (v+u) / (v-u) = 2 / 1, or, v = 3u.